OpenStudy (anonymous):

If because x is differentiable at 2 doesn't necessarily make it continuous at 0, right?

OpenStudy (anonymous):

oh yes it does

OpenStudy (anonymous):

\[\text{differentiable}\rightarrow \text{continuous}\] but not the converse

OpenStudy (anonymous):

But we don't know if its differentiable at 0..

OpenStudy (anonymous):

the fact that \[\lim-{h\to 0}\frac{f(2+h)-f(2)}{h}\] exists implies that \[\lim_{h \to 0} f(2+h)-f(2)=0\] (otherwise the limit of the difference quotient would not exist, and this implies \[\lim_{h\to 0}f(2+h)=f(2)\] means f is continuous at 2

OpenStudy (anonymous):

differentiable at 2 means the limit \[\lim_{h\to 0}\frac{f(2+h)-f(2)}{h}\]exists

OpenStudy (anonymous):

Hmm, that makes sense but my teacher just told me its false

OpenStudy (anonymous):

oh no i am sure what your teacher said is "just because a function is continuous does not mean it is differentiable"

OpenStudy (anonymous):

in other words you can have a function continuous at a point, but not differentiable there. however, if it is differentiable at a point, it is continuous there

OpenStudy (anonymous):

It was a true or false question, stating if x is differentiable at 2, then it is continuous at 0. he said the answer is false.

OpenStudy (anonymous):

oh i am sorry. i did not read correctly. of course. if f is continuous at 2, we know nothing about it at 0. for all we know it doesn't even exists at 0

OpenStudy (anonymous):

kind of a weird question, but it is my fault for not reading carefully

OpenStudy (jamesj):

That's correct. For example, the function f(x) = 1, x > 0 = 0, x =< 0 is differentiable at x=2, but not continuous at x=0.

OpenStudy (anonymous):

Okay, thanks!

OpenStudy (anonymous):

Perhaps you guys can help me with another question i was unsure of.. "Find all of the horizontal asymptotes for y=\[\sqrt{9x^+5}\]/x-1

OpenStudy (anonymous):

\[\sqrt{9x^2+5}\]*